OpenStudy (anonymous):

if a number is subtracted from 91, the principal square root of this difference is equal to the number decreased by 1. What's the number?

OpenStudy (lgbasallote):

a number is subtracted from 91 so this is 91 - x the principal square root of this difference that means positive square root \[+ \sqrt{91 - x}\] equal to the number decreased by 1 so \[+\sqrt{91-x} = x - 1\] does that give you a clue?

OpenStudy (anonymous):

@lgbasallote Might be better to say: \[|\sqrt{91-x}|\]

OpenStudy (lgbasallote):

hmm good point

OpenStudy (lgbasallote):

but isnt that \[\\pm \sqrt{91-x}?\]

OpenStudy (lgbasallote):

\[\pm \sqrt{91-x}*\]

OpenStudy (lgbasallote):

for example \[|x| \implies \pm x\]

OpenStudy (anonymous):

Yeah, that's true. Guess it's just better to say "positive root". Some may get + beforehand as only half of an equation.

OpenStudy (anonymous):

It is 10 \[91-10=81=9^2\]